# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v1

@@ 3.1 Economic Modeling
Government
The government aims to foster economic growth, maintain social equity, and enhance societal welfare through three fiscal tools: debt $B_{t}\in R^{+}$, taxation policy $T_{t}(i_{t},a_{t}):R^{+}\times R^{+}\rightarrow R^{+}$, and spending $G_{t}\in R^{+}$.
$\max_{B_{t},T_{t},G_{t}}\quad\mathbb{E}_{0}\sum_{t=0}^{T_{max}}\left[\alpha_{1}Y_{t}+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}\sum_{i}^{N}u_{t}^{i}\right],$
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are weight parameters reflecting the government’s preference for different objectives. Economic output (e.g., GDP) can be modeled using the Cobb–Douglas production function $Y_{t}=F(K_{t},\mathcal{L}_{t})$, with production capital $K_{t}$ and labor $\mathcal{L}_{t}$. $S(\mathcal{I}_{t},\mathcal{W}_{t})$ denotes the social equity function, where $\mathcal{I}_{t},\mathcal{W}_{t}$ are the income and wealth distributions among households. The sum of households’ utilities $\sum_{i}^{N}u_{t}^{i}$ represents the overall societal welfare.
3.2 Stackelberg Mean Field Game Modeling
In the aforementioned issue, the government (as the leader) formulates policies to guide the decisions of households (as followers), thereby collectively achieving macroeconomic objectives, such as economic growth. Households provide feedback to government policies, creating an asymmetric game with leader-follower dynamics. However, the dynamic interactions among large-scale households extend beyond the standard Stackelberg games. Therefore, this study proposes modeling the problem as a Stackelberg Mean Field Game (Figure 1), based on the following assumptions:
  1. 1.
    Stackelberg assumption: the government (leader) sets policies first, and households (followers) adjust their behavior based on these policies. Both the leader and followers dynamically optimize their strategies, gradually reaching equilibrium.
  2. 2.
    Mean field assumption: households (followers) are assumed to be homogeneous agents. The optimal policies of the leader and followers depend on the overall distribution or average behavior of the followers, rather than a specific one. For example, both governments and households should consider overall market information when making decisions in reality.
Following this, the optimal macroeconomic policy issues can be rewritten as the following game proceeding: For any given time $t\in\{0,\ldots,T_{max}\}$, the leader selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader policy $\pi_{t}^{l}(\cdot|s_{t}^{l})$, given the state $s_{t}^{l}$. Subsequently, the representative follower, located in state $s_{t}^{f}\in\mathcal{S}^{f}$, chooses an action $a_{t}^{f}\in\mathcal{A}^{f}$ according to the policy $\pi_{t}^{f}(\cdot|s_{t}^{f},a^{l}_{t})$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T-1}$ and $\{\pi_{t}^{f}\}_{t=0}^{T-1}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
Followers’ Mean Field Game
Given the leader’s policy $\pi^{l}$, the SMFG is simplified into a mean field game for the followers. At time $t$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s_{t}^{f}$ and action $a_{t}^{f}$ determine the followers’ state-action distribution as $L^{f}_{t}(s_{t}^{f},a_{t}^{f})=\mathbb{P}(s_{t}^{f},a_{t}^{f})$. The representative follower then receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and transitions to the next state $s_{t+1}^{f}\sim P(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L^{f})$ $=\mathbb{E}^{f}\left[\sum_{t=1}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$
$=\mathbb{E}^{f}\left[\sum_{t=1}^{T}u(c_{t},h_{t})\right],$
where $\text{where }c_{t}$ $=W_{t}e_{t}h_{t}+r_{t}a_{t}-T_{t}(i_{t},a_{t})-a_{t+1},$
where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\sim\mu_{0}^{f}$, $s_{t+1}^{f}\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and $a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for each $t$. For households in economic problems, the follower’s state $s^{f}_{t}$ incorporates asset $a_{t}$ and education level $e_{t}$, and the follower’s actions $a^{f}_{t}$ include next savings $a_{t+1}$ and working hours $h_{t}$. The leader’s action $a^{l}_{t}$ denotes tax function $T_{t}(i_{t},a_{t})$, and followers’ distribution $L^{f}_{t}$ includes wage rate $W_{t}$ and interest rate $r_{t}$. Given the leader’s policy $\pi^{l}$, the best response operator of followers is defined as
${\pi^{f}}^{*}\in\mathrm{BR}^{f}(\pi^{l},L^{f})=\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
Stackelberg Leader’s Game
At time $t$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and transition to a new state $s_{t+1}^{l}$ according to the transition probabilities $P(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$. The leader aims to find the best policy $\pi^{l}$ to maximize the total expected rewards:
$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})],$
$r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})=\alpha_{1}Y_{t}+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}\sum_{i}^{N}u_{t}^{i}$
$=\alpha_{1}F(K_{t},\mathcal{L}_{t})+\alpha_{2}S(\mathcal{I}_{t},\mathcal{W}_{t})+\alpha_{3}U(T_{t},W_{t},r_{t})$
where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim P(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T-1\}$. For the government in economic problems, the leader’s state $s^{l}_{t}$ is characterized by production capital $K_{t}$, and the leader’s actions $a^{l}_{t}$ encompass government spending $G_{t}$ and a taxation policy $T_{t}(\cdot)$. The followers’ state-action distribution $L^{f}_{t}$ includes aggregate labor $\mathcal{L}_{t}$, income distribution $\mathcal{I}_{t}$, wealth distribution $\mathcal{W}_{t}$, wage rate $W_{t}$, and interest rate $r_{t}$, these all belong to market information. $U(\cdot)$ represents a measurable function for $T_{t},W_{t},r_{t}$. Therefore, we can define the leader’s best response operator given $\pi^{f},L^{f}$:
${\pi^{l}}^{*}\in\text{BR}^{l}(\pi^{f},L^{f}):=\mathop{\arg\max}_{\pi^{l}}J^{l}(\pi^{l},\pi^{f},L^{f}).$
We then define the transition operator $\mathcal{F}$ as $L^{f}=\mathcal{F}(\pi^{l},\pi^{f},L^{f}_{0},\mu_{0})$ if the evolution of population distribution $L$ satisfies the Mckean-Vlasov equation (1).
$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=$ $\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}|s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})$ (1)
$\quad P(s_{t+1}^{f}|s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}|s_{t+1}^{f}).$
$\mu_{t+1}^{l}(s_{t+1}^{l})=$ $\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}|s_{t}^{l})P(s_{t+1}^{l}|s_{t}^{l},a_{t}^{l}).$
Definition 3.1.
The Stackelberg mean field equilibrium is a tuple $({\pi^{l}}^{*},{\pi^{f}}^{*},{L^{f}}^{*})$, which satisfies the following conditions:
${\pi^{l}}^{*}\in\mathrm{BR}^{l}({\pi^{f}}^{*},{L^{f}}^{*}),{\pi^{f}}^{*}\in\mathrm{BR}^{f}({\pi^{l}}^{*},{L^{f}}^{*}),$
${L^{f}}^{*}=\mathcal{F}({\pi^{l}}^{*},{\pi^{f}}^{*},L^{f}_{0},\mu_{0}).$
4 Methods
Solving the optimal macroeconomic policy is very complex in the real world. Reliance solely on mathematical models for simulation and problem-solving inevitably creates a gap with actual scenarios, thus incorporating real historical data is essential. Nonetheless, a singular focus on historical data for the empirical analysis of economic policies inadequately captures the complex and asymmetric dynamics of interactions between the government and households. Therefore, our solution method is divided into two steps: firstly, we pre-train the follower’s policy based on real data using behavior cloning (BC). Secondly, this paper proposes a model-free Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to solve the Stackelberg Mean Field Games (SMFGs) between the government and large-scale households.
4.1 Pretain by Behavior Cloning
Based on real historical data, we fetch a large number of followers’ state-action pairs $\{s^{f},a^{f}\}$ from a real-data buffer $\mathcal{D}_{real}$ for behavior cloning. For different settings of network structures, we have chosen two types of loss: when the neural network outputs a probability distribution of actions, we use the negative log-likelihood loss (NLL loss); when the neural network outputs action values, we employ the mean square error loss (MSE loss). Our goal is to find the optimal parameters $\theta$ as the follower’s policy network $\pi_{\theta}$ initialization, thereby minimizing the loss to its lowest convergence.
$\min_{\theta}\mathcal{L}_{NLL}$ $=-\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\log{\pi_{\theta}(a^{f}\mid s^{f})},$
$\min_{\theta}\mathcal{L}_{MSE}$ $=\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\left(a^{f}-a\right)^{2}|_{a=\pi_{\theta}(s^{f})}.$
4.2 Stackelberg Mean Field Reinforcement Learning
To solve SMFGs in sequential decision-making, we design a Stackelberg mean field reinforcement learning algorithm based on centralized training with decentralized execution. Consider a game with a leader and $N$ homogeneous follower agents. The leader makes decisions first, and all followers make simultaneous decisions after observing the leader’s actions. The leader aims to find a policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$ to optimize his long-term objective, which is often challenging and requires the cooperation of all followers. Each follower agent seeks to find a policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$ to optimize long-term utility after observing the current leader’s action $a^{l}_{t}=\pi^{l}(s^{l}_{t})$. The joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next observation $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$. Here the experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording experiences of all agents. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})$ is a standard centralized action-value function that takes as input the joint state $\mathbf{s}_{t}$ and joint action $\mathbf{a}_{t}$ and outputs the Q-value for agent $i$. However, due to numerous follower agents, the dimensions of both the joint state and joint action increase significantly with $N$, making the Q-function infeasible to learn. We equate the centralized action-value function $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})$ to $\tilde{Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$ based on mean field theory. The specific theoretical description is shown in the Appendix D.
4.3 Implementation
Furthermore, we consider deterministic policies for both a leader and its followers in a multi-agent system. Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$).
For the leader agent
The policy network $\pi_{\theta^{l}}$ of the leader agent, i.e. the actor, is trained by the sampled policy gradient (Silver et al., 2014):
@@ 4.3 Implementation
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right]$
In the subsequent step, we employ a neural network with parameters $\phi^{l}$ to approximate the leader agent’s action-value function, denoted as $\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ or $\tilde{Q}^{l}$ for brevity. It is updated by minimizing a loss function:
$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$ (2)
@@
where $y^{l}_{t}$ represents the target Q value, computed using the parameters $\phi^{l}_{-}$, the next population distribution $L^{f}_{t+1}$ is based on the Mckean-Vlasov equation (1), and $\gamma$ is discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=$
$\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
For the representative follower agent
The policy network $\pi_{\theta^{f}}$ is trained by the policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}$
$\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
Similar to Mean Field Approximation (Yang et al., 2018), the standard centralized action-value function $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ is approximated as $\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$ with parameters $\phi^{f}$, which is updated by minimizing the MSE loss function:
$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}$
$y^{f}_{t}=$ $r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
@@
where the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$, the next joint followers’ action $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},...,a^{fN}_{t+1}\}$, $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$. Differentiating the loss function $\mathcal{L}(\phi^{f})$ yields the gradient utilized for training:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)$
$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}$
The full algorithm (1) is described in the Appendix.

### v2

@@ 3.2 Dynamic Stackelberg Mean Field Games
where the followers’ state-action distribution $L^{f}$ satisfies the following Mckean-Vlasov equation:
$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l}).$ (3)
In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic Stackelberg mean field game. This modeling approach aims to learn the optimal policy of the leader, namely the social planner, while considering the followers’ best response over discrete timesteps (Definition 3.2).
3.3 Metrics for Dynamic SMFGs
We will utilize the following three metrics to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the followers’ policy from the best response to the leader’s policy; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
Leader’s Payoff
We define the leader’s payoff using the long-term expected rewards of the leader’s policy $\pi^{l}$ over discrete timesteps, as detailed in Equation 2.
Exploitability
Exploitability is a widely used metric for evaluating convergence and measuring the distance between the representative follower’s policy $\pi^{f}$ and the best response ${\pi^{f}}^{*}$ for the leader’s policy $\pi^{l}$. Formally, it quantifies the extent to which a single follower can improve their payoff by deviating from the population’s behavior to adopt a different policy. Given the leader’s policy $\pi^{l}$ and the followers’ state-action distribution $L^{f}$, the exploitability $\mathcal{E}(\pi^{f};\pi^{l})$ of the follower’s policy $\pi^{f}$ is defined by comparing the value functions induced by $\pi^{f}$ and its best response $\pi^{f*}$:
$\mathcal{E}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L^{f})-J^{f}(\pi^{l},{\pi^{f}},L^{f}).$
Social Optimum and Social Welfare
In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 40]. Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions of the $N$ followers in Equation 1 when $N\to\infty$:
$\mathcal{SW}(\pi^{l},\pi^{f})=\sum^{N}J^{f}\left(\pi^{l},\pi^{f},L^{f}\right).$
4 Stackelberg Mean Field Reinforcement Learning
However, solving for the leader’s optimal policy in dynamic SMFGs, as defined in Definition 3.2, presents several challenges, such as vast state and action spaces, strategic dependencies and information asymmetry between the leader and followers, and potentially conflicting objectives including the leader’s payoff, followers’ utilities, and social welfare. These complexities make devising a model-based algorithm with theoretical proof extremely difficult. Previous works have attempted simplifications using the myopic followers setting [63] or linear-quadratic MFGs [6], but this approach fails to align with macroeconomic policy-making issues. Consequently, we have designed the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to address the above limitations. This is a model-free approach that operates independently of environmental transition information.
Figure 2: The architecture of SMFRL algorithm.
Considering the dynamic SMFGs with a leader and $N$ homogeneous follower agents, at any time $t\in\{0,\ldots,T-1\}$, the joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next state $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$. The experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording the experiences of all agents. Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$). The SMFRL algorithm incorporates several settings (shown in Figure 2) to address the aforementioned challenges:
Mean Field Approximation
To mitigate the problem of non-stationarity, we consider a central critic Q function $Q(\mathbf{s}_{t},\mathbf{a}_{t}):\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}\times\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}\to\mathbb{R}$ for each agent. However, due to the high dimensionality of the joint state and action, the standard Q network is infeasible to learn. Thus, the SMFRL algorithm employs a mean field approximation, characterizing the followers’ state-action pairs $(\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t})$ by using the followers’ population distribution $L^{f}_{t}$, which can be presented through various metrics such as mean, variance, probability density functions, skewness, and kurtosis. The theoretical proof is detailed in Appendix F. This approach yields $Q^{l}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ and $Q^{f}(s^{f}_{t},a^{l}_{t},a^{f}_{t},L^{f}_{t})$ for evaluating the leader and follower agents respectively. The mean field approximation effectively reduces the dimensionality of the inputs to the Q function, thereby facilitating the learning of the critic Q network and ensuring the Q function can evaluate leader and follower policies based on the overall follower distribution.
Leader’s Update for Optimal Policy
In SMFRL, the leader’s policy $\pi_{\theta^{l}}$, is optimized using a deterministic policy gradient approach [54]. The policy gradient for the leader can be estimated as:
@@ 4 Stackelberg Mean Field Reinforcement Learning
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right],$
where ${Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$, the leader’s action-value function, is approximated by a neural network parameterized by $\phi^{l}$. This network is periodically updated to minimize the loss:
$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
where the target value $y^{l}_{t}$ is given by:
$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$
@@
and $\phi^{l}_{-}$ represents the parameters of the target Q network, $L^{f}_{t+1}$ is determined based on Mckean-Vlasov equation (3), and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right].$
Followers’ Update for Best Response
To enable follower agents to effectively learn the best response policy $\pi^{f*}$ to the leader’s policy $\pi^{l}$, we employ multiple updates for followers’ shared policy and Q networks in the SMFRL algorithm. Each update cycle refines the policy network $\pi_{\theta^{f}}$ for the followers using the policy gradient method:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f}_{t})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
where the action-value function ${Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$, parameterized by $\phi^{f}$, is periodically updated by minimizing the mean squared error (MSE) loss:
$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left[\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$ (4)
$y^{f}_{t}=$ $r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
@@
with the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, and the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$. The joint actions of followers $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},\ldots,a^{fN}_{t+1}\}$ are determined with each action $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
The pseudocode for the SMFRL algorithm (1) is presented in Appendix A.
Table 1: Evaluation of SMFRL and five baselines on 3 metrics for dynamic Stackelberg mean field games with different followers’ numbers.
Metrics Followers’ Optimal SMFRL MF-MARL MADDPG IDDPG DDPG U.S. Policy
Number Value +Real +Real
Leader’s Payoff N=100 \ 3387.36 -76.82 -282.62 -701.61 2840.07 1718.08
N=1000 \ 3411.16 -668.03 121.18 -767.53 2874.19 -796.4
Exploitability N=100 0. 0.000107 0.002190 0.007908 0.001333 0.009885 0.032490
N=1000 0. 0.000158 0.003481 0.035613 0.000565 0.003002 0.010522
Social Welfare N=100 100 96.87 82.02 55.74 83.09 94.5 79.23
N=1000 1000 968.94 834.93 645.38 874.53 943.82 334.79